### v1

@@ 3.2 AdaCUR algorithm for accuracy
(ii) Low-cost minor modifications
@@
The low-cost minor modifications involve trying to enlarge the index set using the sketch of the residual error matrix $E_{s}$ (lines 1010101011111111), reorder them by importance (lines 1212121213131313), remove the indices that are not important (lines 1414141417171717) and recompute the relative error (lines 1818181819191919). More specifically, in line $10$, randomized pivoting is used (Algorithm 1) where the sketch of the residual matrix, $E_{s}$ is used to obtain an additional set of $s$ column indices and $A(t_{j})([m]-I,J_{1})^{T}$ is used to obtain an additional set of $s$ row indices. We append the extra set of $s$ indices to the original set of indices $I$ and $J$ in line $11$ and order the indices in terms of their importance in lines 1212121213131313. Here, Gu-Eisenstat’s strong rank-revealing QR factorization [27] is used, but we can use other strong rank-revealing algorithms or more expensive algorithms with strong theoretical guarantee such as [14], because we are working with a small $O(r)\times r$ matrix $A(t_{j})(I\cup I_{0},J)$. In line $14$, we approximate the $(\epsilon/\sqrt{n})$-rank of $A(t_{j})(I\cup I_{0},J)$ by looking at the diagonal entries of the upper triangular factor in sRRQR as they give an excellent approximation to the singular values of the original matrix. [^10] Instead of approximating the $(\epsilon/\sqrt{n})$-rank of $A(t_{j})(I\cup I_{0},J)$, we can set the rank tolerance slightly smaller, for example, $0.5\epsilon/\sqrt{n}$, to account for the approximation error and decrease the chance of recomputing the indices from scratch (lines 2121212122222222) by keeping slightly more indices than necessary. In lines 1515151517171717, we truncate based on the rank computed in line $14$. The ordering of the indices is important here. We recompute the relative error in lines 1818181819191919, [^11] For theoretical guarantee, the Gaussian embedding $\Gamma_{s}$ needs to be independent from the modified indices [26], however the extra set of indices is dependent on $\Gamma_{s}$ as we apply pivoting on the sketched residual in line $10$. Nonetheless, the dependency is rather weak so $\Gamma_{s}$ can be reused without losing too much accuracy. and if the relative error $E$ is smaller than the tolerance, we store the new set of columns, rows of $A(t_{j})$ and their intersection in line $23$ and continue to the next parameter value. If the relative error exceeds the tolerance, there are two recourses. AdaCUR outlines the first option in lines 2121212122222222 where we recompute the indices altogether from scratch. In line $21$, we perform both randomized rank estimation and randomized pivoting for two reasons. First, randomized rank estimation is extremely cheap when we have the row sketch already and second, when there is a sudden large change in rank, we can use randomized rank estimation to adjust the rank efficiently; note that when minor modifications did not help in lines 1010101017171717, the matrix may have changed drastically. An alternative approach, not included in AdaCUR, is to increase the error sample size $s$ gradually by incrementing its value, say $2s,4s,$ and so on, and updating the sketch accordingly by appending to it. As $s$ increases, the number of additional important indices obtained from randomized pivoting in line $10$ also increases, which will eventually reduce the relative error to less than the tolerance $\epsilon$.
@@
The value of $s$ can be important for two reasons: error estimation and minor modifications. If we set $s$ to be larger then we obtain higher accuracy in our error estimation at the cost of more computation, but more importantly, we obtain a larger set of extra indices for the minor modification. This can help the algorithm avoid the if statement in line $20$, which recomputes the indices from scratch; see also Section 4.3. We can also create a version with increasing $s$ if minor modifications fail to decrease the relative error below the tolerance, as described in the final part of the previous paragraph.
@@
(12) $\mathcal{O}\left(\sum\limits_{j\in Q_{1}}\left(r_{j}T_{A(t_{j})}+(m+n)r_{j}^{2}+nr_{j}(r_{j}+p)\right)+\sum\limits_{j\in Q_{2}}\left(sT_{A(t_{j})}+(m+n)r_{j}s+nsp\right)\right)$
@@
It turns out that in many applications, the cardinality of the set $Q_{1}$ is small when we choose the oversampling parameter $p$ and the error sample size $s$ appropriately, making the algorithm closer to $\mathcal{O}(T_{A(t_{j})})$ for the $j$th parameter value; see Sections 4.2 and 4.3. This makes AdaCUR usually faster than other existing methods such as dynamical low-rank approximation [41], and randomized SVD and generalized Nyström method [37]; see Section 4.5.
@@ 3.3 FastAdaCUR algorithm for speed
The main distinction between AdaCUR and FastAdaCUR lies in the existence of error estimation. In AdaCUR, an error estimate is computed for each parameter value, which ensures that the row and the column indices are guaranteed to be good with high probability. This is absent in FastAdaCUR. Consequently, if there is a sudden change in rank or the previous set of indices are unimportant for the current parameter value, FastAdaCUR may deliver poor results. To mitigate this disadvantage, FastAdaCUR incorporates a buffer space (i.e., additional indices in $I,J$) as input to assist with detecting changes in rank. FastAdaCUR involves many heuristics, so we break it down to discuss and justify each part of the algorithm. We divide FastAdaCUR into three parts: (i) rank estimation using the core matrix (lines 888811111111), (ii) rank decrease via truncation (lines 1212121214141414) and (iii) rank increase via oversampling (lines 1515151518181818). The initial set of indices are constructed from scratch as in Section 3.1, with the indices for the buffer space and oversampling obtained using the oversampling algorithm from Algorithm 4.
(i) Rank estimation using the core matrix
@@
Upon entering the for loop in line $7$, we first compute the core matrix $U=A(t_{j})(I,J)\in\mathbb{R}^{(r+b+p)\times(r+b)}$ (line $8$) for the current parameter value, which includes extra indices from the buffer space and oversampling. As FastAdaCUR prioritizes efficiency over accuracy, the core matrix is a sensible surrogate to approximate the singular values of $A(t_{j})$ without looking at the entire matrix. The extra indices from the buffer space and oversampling will assist with the approximation as the buffer space allows the core matrix to approximate more singular values, helping to detect a potential $(\epsilon/\sqrt{n})$-rank change. The additional set of row indices from oversampling improves the accuracy of the estimated singular values of $A(t_{j})$ (line $10$) by the Courant-Fischer min-max theorem, which is used for approximating the relative $(\epsilon/\sqrt{n})$-rank in line $11$. In addition, oversampling increases the singular values of the core matrix, thereby allowing FastAdaCUR to increase the rank when appropriate, which helps the algorithm to reduce the error; however, we see in the experiments (Section 4.2) that the role that oversampling plays in FastAdaCUR is rather complicated. We use sRRQR to order the columns and rows of the core matrix by importance in lines $9$ and $10$, and subsequently estimate the $(\epsilon/\sqrt{n})$-rank of the core matrix $U$ using the diagonal entries of the upper triangular factor in sRRQR in line $11$. We use the $(\epsilon/\sqrt{n})$-rank to aim for a tolerance of $\epsilon$ in the low-rank approximation. If the computed rank, $r_{0}$ is smaller than the previous rank, $r$, then we decrease the index set via truncation. Otherwise, we increase the index set via oversampling. The details are discussed below.
(ii) Rank decrease via truncation
@@
(iii) Rank increase via oversampling
If the rank has increased, i.e. $r_{0}>r$, we refill the extra indices by the amount the rank has increased by, i.e., $r_{0}-r$. Unlike in AdaCUR, which we have the sketched residual matrix, we do not have a good proxy to obtain a good set of extra indices. Therefore, we use the already-selected columns and rows to obtain an extra set of column and row indices. [^13] The only recourse to obtaining a reliable set of extra indices is to view the entire matrix; see [43, § 17.5] for a related discussion. However, this puts us in a similar setting as AdaCUR, which we recommend when ensuring accuracy is the priority rather than speed. We use the oversampling algorithm (Algorithm 4) to achieve this as it only requires the selected rows and columns as input. Adding extra indices using oversampling has been suggested before in [17] to make small rank increases in the context of dynamical low-rank approximation. We get the extra indices from oversampling in lines $16$ and $17$ and append to the original set of indices in line $18$. At the end of this procedure, the algorithm ensures $|I|=r_{0}+b+p$ and $|J|=r_{0}+b$.
@@
Let $r_{j}$ be the size of the index set for parameter $t_{j}$, i.e., the rank of the CUR decomposition for $A(t_{j})$. The dominant cost $\mathcal{O}\left(r_{1}T_{A(t_{1})}\right)$ in FastAdaCUR comes from line $1$ for computing the initial set of indices for $A(t_{1})$ excluding the buffer space and oversampling. Besides the computation of initial indices, the complexity is at most linear in $m$ and $n$. For $j=2,...,q$, the complexity of FastAdaCUR for the $j$th parameter value is $\mathcal{O}((m+n)(r_{j}+p+b)^{2})$ if the rank has increased from $t_{j-1}$ to $t_{j}$. The cost involves the usage of oversampling to replenish extra indices. If the rank has either decreased or stayed the same then the complexity is $\mathcal{O}((r_{j}+p+b)(r_{j}+b)^{2})$ for using sRRQR. Hence, excluding the first line, FastAdaCUR runs at most linearly in $m$ and $n$ for each parameter value, making it remarkably fast as demonstrated in Section 4.5.
@@ 4 Numerical Illustration
In the following sections, we perform different numerical simulations with varying input parameters to test AdaCUR and FastAdaCUR. We test
@@
    the speed of our algorithms against existing methods in [37, 41],
using various problems in their respective sections. It is worth pointing out that in the first three examples $\min_{j}\left\lVert A(t_{j})-A(t_{j-1})\right\rVert_{2}>0.03$ for all $j$, indicating that the performance of AdaCUR and FastAdaCUR is not explained by the fact that the matrices are undergoing small perturbations as we iterate through the parameters.
In the experiments, we measure how many times AdaCUR need to enter some of the if statements, as the heavier computations are usually done inside them. This happens when the current set of indices do not meet the given tolerance in AdaCUR. In FastAdaCUR, besides the computation of initial set of indices, the heaviest computation is done when the rank has increased, i.e. oversampling in lines $16$ and $17$ is invoked. However, this is heavily dependent on the given problem, so we do not track this. We define the following
@@
In the plots that depict the $(\epsilon/\sqrt{n})$-rank of the low-rank approximation, a large dot is used to indicate the parameter value for which the indices were recomputed from scratch in AdaCUR. The number of large dots is equal to $h_{2}$. For reference in FastAdaCUR, a large dot is used to indicate the parameter value where oversampling was applied when the rank increased.
@@ 4.1 Varying the tolerance $\epsilon$
In this section, we test the accuracy of our algorithms by varying the tolerance parameter $\epsilon$. We use the synthetic example given in [9], which tests the robustness against small singular values. The example is given by
(14) $A(t)=e^{tW_{1}}e^{t}De^{tW_{2}},t\in[0,1]$
where $D\in\mathbb{R}^{n\times n}$ is diagonal with entries $d_{jj}=2^{-j}$ and $W_{1},W_{2}\in\mathbb{R}^{n\times n}$ are randomly generated skew-symmetric matrices. The singular values of $A(t)$ are $e^{t}2^{-j}$ for $j=1,2,...,n$. We take $n=500$ and take $101$ equispaced points in the interval $[0,1]$ for $t$. The matrix exponential was computed using the $\mathtt{expm}$ command in MATLAB.
The results are depicted in Figure 1 and Table 1. In Figures 1a and 1b, we run AdaCUR with varying tolerance $\epsilon$ on the synthetic problem (14). We set the oversampling parameter $p$ to $5$ and the error sample size $s$ to $5$ in this experiment. We first observe in Figure 1a that the relative error for AdaCUR meets the tolerance $\epsilon$ within a small constant factor. This is expected, given that the randomized methods in the algorithm typically yield errors within a small constant factor of the optimal solution; therefore one can simply take a smaller $\epsilon$ to ensure the actual error is bounded by the desired accuracy. Furthermore, the performance of AdaCUR is reflected in Figure 1b where the algorithm attempts to efficiently find the low-rank approximation of the parameter-dependent matrix by matching the $(\epsilon/\sqrt{n})$-rank at each parameter value. Regarding the algorithm’s complexity, as shown in Table 1, we observe that $h_{2}$ increases as the tolerance $\epsilon$ increases. This increase is anticipated since higher accuracy requirement generally equates to more heavy computations.
In Figures 1c and 1d, we run FastAdaCUR with varying tolerance $\epsilon$ on the synthetic problem (14). We set the oversampling parameter $p$ to $5$ and the buffer space $b$ to $5$. We notice in Figure 1c that FastAdaCUR performs well by meeting the tolerance $\epsilon$ within a small constant factor. This performance is also reflected in Figure 1d, where the algorithm tries to match the $(\epsilon/\sqrt{n})$-rank at each parameter value. However, as shown in the numerical examples that follow, FastAdaCUR must be used with caution because as the problem becomes more difficult, the error may continue to grow as we iterate through the parameter values.
@@
Minor mod. only $h_{1}$ $16$ $19$ $19$ $17$
Recomput. Indices $h_{2}$ $22$ $30$ $34$ $41$

### v2

@@ 3.2.1 Initial error estimation
3.2.2 Low-cost minor modifications
@@
The low-cost modifications involve trying to enlarge the index set using the sketch of the residual error matrix $E_{s}$ (lines 1010101011111111), reordering them by importance (lines 1212121213131313), removing unimportant indices (lines 1414141417171717), and recomputing the relative error (lines 1818181819191919). More specifically, in line $10$, randomized pivoting is used on the sketch of the residual matrix, $E_{s}$ as in Algorithm 4 to obtain an additional set of $s$ row and column indices. We append the extra set of $s$ indices to the original sets of indices $I$ and $J$ in line $11$, and order the indices in terms of their importance in lines 1212121213131313. Here, Gu-Eisenstat’s strong rank-revealing QR factorization [27] is used, but we can use other strong rank-revealing algorithms or more expensive algorithms with strong theoretical guarantees such as [14, 49], because we are working with a small $O(r)\times r$ matrix $A(t_{j})(I\cup I_{0},J)$. In line $14$, we approximate the $(\epsilon/\sqrt{n})$-rank of $A(t_{j})(I\cup I_{0},J)$ by looking at the diagonal entries of the upper triangular factor in sRRQR as they give an excellent approximation to the singular values of the original matrix. [^9] Instead of approximating the $(\epsilon/\sqrt{n})$-rank of $A(t_{j})(I\cup I_{0},J)$, we can set the rank tolerance slightly smaller, for example, $0.5\epsilon/\sqrt{n}$, to account for the approximation error and decrease the chance of recomputing the indices from scratch (lines 2121212122222222) by keeping slightly more indices than necessary. In lines 1515151517171717, we truncate based on the rank computed in line $14$. The ordering of the indices is important here. We recompute the relative error in lines 1818181819191919, [^10] For theoretical guarantee, the Gaussian embedding $\Gamma_{s}$ needs to be independent from the modified indices [26], however the extra set of indices is dependent on $\Gamma_{s}$ as we apply pivoting on the sketched residual in line $10$. Nonetheless, the dependency is rather weak so $\Gamma_{s}$ can be reused without losing too much accuracy. and if the relative error $E$ is smaller than the tolerance, we store the new set of columns, rows of $A(t_{j})$ and their intersection in line $23$ and continue to the next parameter value. If the relative error exceeds the tolerance, there are two recourses. AdaCUR outlines the first option in lines 2121212122222222 where we recompute the indices altogether from scratch. In line $21$, we perform both randomized rank estimation and randomized pivoting for two reasons. First, randomized rank estimation is extremely cheap when we have the row sketch already and second, when there is a sudden large change in rank, we can use randomized rank estimation to adjust the rank efficiently; note that when minor modifications did not help in lines 1010101017171717, the matrix may have changed drastically.
An alternative to recomputing the indices
An alternative approach, not included in AdaCUR, is to increase the error sample size $s$ gradually by incrementing its value, say $2s,4s,$ and so on, and updating the sketch accordingly by appending to it. As $s$ increases, the number of additional important indices obtained from randomized pivoting in line $10$ also increases, which will eventually reduce the relative error to less than the tolerance $\epsilon$.
@@
The value of $s$ can be important for two reasons: error estimation and minor modifications. If we set $s$ to be larger, we obtain higher accuracy in our error estimation at the cost of more computation. However, we obtain a larger set of extra indices for the minor modification. This can help the algorithm avoid the if statement in line $20$, which recomputes the indices from scratch; see also Section 4.3. We can also create a version with increasing $s$ if minor modifications fail to decrease the relative error below the tolerance, as described in the final part of the previous paragraph.
@@
(14) $\mathcal{O}\Big{(}\sum\limits_{j\in Q_{1}}\left(r_{j}T_{A(t_{j})}+(m+n)r_{j}^{2}+nr_{j}(r_{j}+p)\right)+\sum\limits_{j\in Q_{2}}\left(sT_{A(t_{j})}+(m+n)r_{j}s+nsp\right)\Big{)}$
@@
It turns out that in many applications, the cardinality of the set $Q_{1}$ is small when we choose the oversampling parameter $p$ and the error sample size $s$ appropriately, making the algorithm closer to $\mathcal{O}(T_{A(t_{j})})$ for the $j$th parameter value; see Sections 4.2 and 4.3. This makes AdaCUR usually faster than other existing methods such as dynamical low-rank approximation [41], and randomized SVD and generalized Nyström method [37]; see Section 4.5.2.
@@ 3.3 FastAdaCUR algorithm for speed
The main distinction between AdaCUR and FastAdaCUR lies in the existence of error estimation. In AdaCUR, an error estimate is computed for each parameter value, which ensures that the row and the column indices are guaranteed to be good with high probability. This is absent in FastAdaCUR. Consequently, if there is a sudden change in rank or the previous set of indices are unimportant for the current parameter value, FastAdaCUR may deliver poor results. To mitigate this disadvantage, FastAdaCUR incorporates a buffer space (i.e., additional indices in $I,J$) as input to assist with detecting changes in rank. FastAdaCUR involves many heuristics, so we break it down to discuss and justify each part of the algorithm. We divide FastAdaCUR into three parts:
  • Section 3.3.1: Rank estimation using the core matrix (lines 888811111111),
  • Section 3.3.2: Rank decrease via truncation (lines 1212121214141414),
  • Section 3.3.3: Rank increase via oversampling (lines 1515151518181818).
The initial set of indices are constructed from scratch as in Section 3.1, with the indices for the buffer space and oversampling obtained using the oversampling algorithm from Algorithm 5.
3.3.1 Rank estimation using the core matrix
@@
Upon entering the for loop in line $7$, we first compute the core matrix $U=A(t_{j})(I,J)\in\mathbb{R}^{(r+b+p)\times(r+b)}$ (line $8$) for the current parameter value, which includes extra indices from the buffer space and oversampling. As FastAdaCUR prioritizes efficiency over accuracy, the core matrix is a sensible surrogate to approximate the singular values of $A(t_{j})$ without looking at the entire matrix. The extra indices from the buffer space and oversampling will assist with the approximation as the buffer space allows the core matrix to approximate more singular values, helping to detect a potential $(\epsilon/\sqrt{n})$-rank change. The additional set of row indices from oversampling improves the accuracy of the estimated singular values of $A(t_{j})$ (line $10$) by the Courant-Fischer min-max theorem, which is used for approximating the relative $(\epsilon/\sqrt{n})$-rank in line $11$. In addition, oversampling increases the singular values of the core matrix, thereby allowing FastAdaCUR to increase the rank when appropriate and reduce the error. However, we see in the experiments (Section 4.2) that the role that oversampling plays in FastAdaCUR is rather complicated. We use sRRQR to order the columns and rows of the core matrix by importance in lines $9$ and $10$, and subsequently estimate the $(\epsilon/\sqrt{n})$-rank of the core matrix $U$ using the diagonal entries of the upper triangular factor in sRRQR in line $11$. We use the $(\epsilon/\sqrt{n})$-rank to aim for a tolerance of $\epsilon$ in the low-rank approximation. If the computed rank, $r_{0}$ is smaller than the previous rank, $r$, then we decrease the index set via truncation. Otherwise, we increase the index set via oversampling. The details are discussed below.
3.3.2 Rank decrease via truncation
@@
3.3.3 Rank increase via oversampling
If the rank has increased, i.e. $r_{0}>r$, we refill the extra indices by the amount the rank has increased by, i.e., $r_{0}-r$. Unlike in AdaCUR, for which we have the sketched residual matrix, we do not have a good proxy to obtain a good set of extra indices. Therefore, we use the already-selected columns and rows to obtain an extra set of column and row indices. [^12] The only recourse to obtaining a reliable set of extra indices is to view the entire matrix; see [43, § 17.5] for a related discussion. However, this puts us in a similar setting as AdaCUR, which we recommend when ensuring accuracy is the priority rather than speed. We use the oversampling algorithm (Algorithm 5) to achieve this as it only requires the selected rows and columns as input. Adding extra indices using oversampling has been suggested before in [17] to make small rank increases in the context of dynamical low-rank approximation. We get the extra indices from oversampling in lines $16$ and $17$ and append to the original set of indices in line $18$. At the end of this procedure, the algorithm ensures $|I|=r_{0}+b+p$ and $|J|=r_{0}+b$.
@@ 3.3.3 Rank increase via oversampling
Let $r_{j}$ be the size of the index set for parameter $t_{j}$, i.e., the rank of the CUR decomposition for $A(t_{j})$. The dominant cost $\mathcal{O}\left(r_{1}T_{A(t_{1})}\right)$ in FastAdaCUR comes from line $1$ for computing the initial set of indices for $A(t_{1})$ excluding the buffer space and oversampling. Besides the computation of initial indices, the complexity is at most linear in $m$ and $n$. For $j=2,...,q$, the complexity of FastAdaCUR for the $j$th parameter value is $\mathcal{O}((m+n)(r_{j}+p+b)^{2})$ if the rank has increased from $t_{j-1}$ to $t_{j}$. The cost involves the usage of oversampling to replenish extra indices. If the rank has either decreased or stayed the same then the complexity is $\mathcal{O}((r_{j}+p+b)(r_{j}+b)^{2})$ for using sRRQR. Hence, excluding the first line, FastAdaCUR runs at most linearly in $m$ and $n$ for each parameter value, making it remarkably fast as demonstrated in Section 4.5.2.
@@ 4 Numerical Illustration
In the following sections, we perform different numerical simulations with varying input parameters to test AdaCUR and FastAdaCUR. These simulations allow us to test the following.
@@
    The speed of our algorithms against existing methods in [37, 41].
It is worth pointing out that in the first three examples $\min_{j}\left\lVert A(t_{j})-A(t_{j-1})\right\rVert_{2}>0.03$ for all $j$, indicating that the performance of AdaCUR and FastAdaCUR is not explained by the fact that the matrices are undergoing small perturbations as we iterate through the parameters.
In the experiments, we measure how many times AdaCUR needs to enter some of the ‘if’ statements, as the heavier computations are usually done inside them. This happens when the current set of indices do not meet the given tolerance in AdaCUR. We define the following
@@
In the plots that depict the $(\epsilon/\sqrt{n})$-rank of the low-rank approximation, a large dot is used to indicate the parameter value for which the indices were recomputed from scratch in AdaCUR. The number of large dots is equal to $h_{2}$. In FastAdaCUR, besides the computation of initial sets of indices, the heaviest computation is done when the rank has increased; i.e. in lines $16$ and $17$ when oversampling is invoked. However, this is heavily dependent on the given problem, so we do not track this. For reference, in FastAdaCUR, a large dot is used to indicate the parameter value where oversampling was applied when the rank increased.
@@ 4.1 Varying the tolerance $\epsilon$
In this section, we test the accuracy of our algorithms by varying the tolerance parameter $\epsilon$. We use the same synthetic example [9] used for Figure 3, which is given by
(13) $A(t)=e^{tW_{1}}e^{t}De^{tW_{2}},t\in[0,1].$
This example tests robustness against small singular values.
The results are depicted in Figure 4 and Table 1. In Figures 4a and 4b, we run AdaCUR with varying tolerance $\epsilon$ on the synthetic problem (13). We set the oversampling parameter $p$ to $5$ and the error sample size $s$ to $5$ in this experiment. We first observe in Figure 4a that the relative error for AdaCUR meets the tolerance $\epsilon$ within a small constant factor. This is expected, given that the randomized methods in the algorithm typically yield errors within a small constant factor of the optimal solution; therefore one can simply take a smaller $\epsilon$ to ensure the actual error is bounded by the desired accuracy. Furthermore, the performance of AdaCUR is reflected in Figure 4b where the algorithm attempts to efficiently find the low-rank approximation of the parameter-dependent matrix by matching the $(\epsilon/\sqrt{n})$-rank at each parameter value. Regarding the algorithm’s complexity, as shown in Table 1, we observe that $h_{1}$ and $h_{2}$ increases as the tolerance $\epsilon$ decreases. This increase is anticipated since a higher accuracy requirement generally equates to heavier computations.
In Figures 4c and 4d, we run FastAdaCUR with varying tolerance $\epsilon$ on the synthetic problem (13). We set the oversampling parameter $p$ to $5$ and the buffer space $b$ to $5$. We notice in Figure 4c that FastAdaCUR performs well by meeting the tolerance $\epsilon$ within a small constant factor. This performance is also reflected in Figure 4d, where the algorithm tries to match the $(\epsilon/\sqrt{n})$-rank at each parameter value. However, as shown in the numerical examples that follow, FastAdaCUR must be used with caution because as the problem becomes more difficult, the error may continue to grow as we iterate through the parameter values.
@@
Minor mod. only $h_{1}$ $6$ $8$ $10$ $11$
Recomput. Indices $h_{2}$ 00 00 00 $1$